Your rich uncle just gave you a gift of $40,000. You are investing this money for 15 years at 7 percent simple interest. How much money will you have at the end of the 15 years?
Found 2 solutions by Boreal, ikleyn:
Answer by Boreal(15235)

I=prt=40000*0.07*15=$42,000 of interest
But the principal is also part of this, so the total will be $82,000

Simple interest compounded (not the question) would be P=Po(1+r)^15
P=40000(1+.07)^15, rule of 72 would say this will double in 10 years, so expect answer to be well over $80,000 but doesn't double twice, so below $160,000
$110,361.26 Compounding matters!

Answer by ikleyn(52782)

.
40000*(1 + 0.07*15) dollars.

Use your calculator.

------------------

Notice that @Boreal misread the problem and gave the solution and the answer not for the simple interest, but for compounded account.